### v1

# Effect of Adaptive Communication Support on Human-AI Collaboration

Abstract
Effective human-AI collaboration requires agents to adopt their roles and levels of support based on human needs, task requirements, and complexity. Traditional human-AI teaming often relies on a pre-determined robot communication scheme, restricting teamwork adaptability in complex tasks. Leveraging strong communication capabilities of Large Language Models (LLMs), we propose a Human-Robot Teaming Framework with Multi-Modal Language feedback (HRT-ML), a framework designed to enhance human-robot interaction by adjusting the frequency and content of language-based feedback. HRT-ML framework includes two core modules: a Coordinator for high-level, low-frequency strategic guidance, and a Manager for task-specific, high-frequency instructions, enabling passive and active interactions with human teammates. To assess the impact of language feedback in collaborative scenarios, we conducted experiments in an enhanced Overcooked-AI game environment with varying levels of task complexity (easy, medium, hard) and feedback frequency (inactive, passive, active, superactive). Our results show that as task complexity increases relative to human capabilities, human teammates exhibited stronger preference towards robotic agents that can offer frequent, proactive support. However, when task complexities exceed the LLM’s capacity, noisy and inaccurate feedback from superactive agents can instead hinder team performance, as it requires human teammates to increase their effort to interpret and respond to the large amount of communications, with limited performance return. Our results offers a general principle for robotic agents to dynamically adjust their levels and frequencies of communications to work seamlessly with human and achieve improved teaming performance.
Introduction
Human-robot collaboration has been extensively studied and applied across diverse scenarios, demonstrating strong potential towards enhanced efficiency and performance (???????). As task complexity increases, agent adaptability becomes increasingly essential for seamless teamwork. Previous work has developed methods and tools for robot to adapt their actions based on inferred human objectives (?), trust level (?), and individual preferences (?). Recent work also begin to incorporate language-based feedback (??) to enable more direct communications and lower user barriers. However, robot communications in these approaches primarily focused on relative simple commands such as “pick up the book” or “move left a little”, which do not reflect the level of human-robot communications required in real-world applications.
Recent advancements in large language models (LLMs) have brought powerful reasoning (????), natural language understanding (??), contextual awareness (?), and generalization capabilities (?), enabling more advanced, prolonged communications (???). These methods has empowered agents to process ambiguous and complex instructions from human (?), engage in more natural and dynamic conversations (??), and learn from a diverse set of inputs (??).
However, even in these LLM-enhanced communications, human teammates continue to play a predominant role in requesting specific tasks and providing suggestions during collaboration (?). Schoenegger et al. (?)’s study suggested that state-of-the-art LLMs can often match or surpass human performance in various domains. Based on these results, we hypothesize that allowing LLM agents to more proactively participate in (?) or even initiate communications with human teammates can enhance teaming performance and efficiency.
To test this hypothesis and systematically evaluate the impact of different forms of language feedback provided by robots on collaboration efficiency and human satisfaction, in this study we develop HRT-ML, a human-robot teaming framework incorporating multi-modal language feedback to support dynamic, context-aware teaming styles. To enable the robot to provide effective language feedback, HRT-ML comprises two main modules: a Coordinator, which manages overall collaboration strategies and delivers low-frequency or passive instructions and feedback, and a Manager, which determines appropriate subtasks based on the coordinated plan at each stage, offering high-frequency instructions. Combining Coordinator and Manager allows the robotic agent to provide different forms of instruction at different frequencies. To investigate how the form and frequency of language feedback influence teaming performance, we performed user studies using four different agent active levels: Inactive, Passive, Active, and Superactive. We find that as the environment becomes more challenging, participants exhibit stronger preference for agents that provide active support, whereas in simpler tasks, frequent agent feedback is often perceived negatively, reducing overall performance and human satisfaction. These findings reveal the importance for the active level of robot language support to dynamically adapt based on task complexity and team capability. Inaccurate or overly-frequent feedback can decrease team efficiency, as participants must take extra time to understand and correct these suggestions.
In summary, our contributions are as follows:
  • developed a human-robot teaming framework incorporating multi-modal language feedback to support dynamic, context-aware teaming styles.
  • performed user studies to determine how forms and frequency of agent language feedback influence teaming performance
  • discovered how the active level of robot language support should adapt based on task complexity and team capability, to best enhance teaming performance
To the best of the authors’ knowledge, this is the first study that systematically explores the effect of different types of LLM-based multimodal language feedback on teaming performance across a wide range of task complexities.
Testbed: Overcooked-AI
To test the influence of language feedback on human-robot collaboration, we chose Overcooked-AI (?), a platform designed to assess multi-agent coordination skills. In this game, players are motivated to collaborate actively to maximize their score by completing orders within a time limit. A score of 60 points was awarded when the correct soup is served. Partial points will be awarded to incomplete or incorrect soups based on the number of missing/incorrect components. To cook a soup, chefs need to finish specific subtasks in sequence according to the recipes (see Fig. 1A), and to finish subtasks, chefs need to move and interact with the environment. This process can be applied to any collaborative setting: first, reasoning through subtasks to achieve the overall goal, then selecting low-level atomic actions to complete each subtask.
Figure 1: (A) Cooking process to complete an order. (B) The designed human-AI collaboration interface (left: game layout, right: communication panel). The red cross represent an example of the intermediate empty counter used to collaborate.
Subtasks
The subtasks in an overcooked environment for a single agent can be broken down into three main parts:
  1. 1.
    Gathering Ingredients: Chefs must first pick up the correct ingredients, such as onions or tomatoes, and place them into the cooking pot according to the recipe requirements.
  2. 2.
    Cooking: Once the ingredients are placed in the pot, the agent has to start cooking. A timer on the pot signals when the soup is ready to be served.
  3. 3.
    Serving: When the soup is ready, the chefs must collect and clean the dish from the dish dispenser, pour the soup into the dish, and deliver it to the serving location.
In multi-agent collaborative scenarios, each agent has a different path cost for completing each subtask, such as picking up an onion versus a tomato. Some subtasks may be unachievable for certain agents. Furthermore, by decomposing a subtask into multiple smaller subtasks that can be performed by multiple agents, the overall time cost can be reduced. For example, in Fig. 1, the green agent might place the onion at the red cross point (5,2), allowing the blue agent to pick it up from there and add it to the pot.
Atomic Action
To finish a specific subtask, such as picking up an onion, the agent must execute a sequence of atomic actions, including movement commands like up, down, left, right, stay, and interact for picking up or placing objects.
Figure 2: Flowing Human-Robot Teaming Framework (HRT-ML). It contains two modules: the Coordinator and the Manager. The Coordinator considers human preferences, formulates overall strategies, and provides low-frequency feedback to guide the collaboration. The Manager processes these strategies along with the state information to generate high-frequency subtask instructions for both human players and greedy planners, facilitating efficient task coordination and execution.
Human-Robot Teaming Framework with Multi-Modal Language Feedback
To provide adaptive language feedback in human-robot collaboration, HRT-ML includes two core components: the Coordinator and the Manager. The Coordinator leads high-level strategy discussions with humans, generating a final coordination plan and offering low-frequency feedback. In contrast, the Manager handles detailed subtask allocation and provides high-frequency feedback to guide human players. Humans and other agents will receive the feedback and execute low-level actions.
Coordinator
The Coordinator is responsible for designing overall collaboration strategies and discussing them with the human partner. To support this, we employed a structured chain-of-thought approach to make the suggestions and discussions more effective. It begins by retrieving relevant information, including the layout, the overall collaboration goal, rules, and the agent’s state formatted as a text description (Figure 2) using a prompt template, and then queries GPT-4o with it. It first analyzes the possible subtasks required to achieve the overall goal, then evaluates the difficulty of each subtask for each agent, and generates a plan that maximizes collaboration efficiency. For example, based on the environment (Fig. 1B), our Coordinator suggests: Given the layout, the human (blue agent) should focus on picking up the tomato and serving the dish, while I will handle the onion and place the dish at location (5, 2) or (6, 2). We note that the coordination querying is made conversational, allowing humans to continuously revise the plan through continuous discussion after reviewing the plan. During these coordination discussions, the game remains paused until the human chooses to end the conversation. The conversation can be initiated proactively by the Coordinator at a low frequency or by the human. When a human request is made, we incorporate their suggestions into the coordination plan prompt for initially querying GPT-4, explicitly asking it to evaluate the feasibility of these preferences and propose a corresponding coordination plan.
Manager
The Manager assigns subtasks at each stage, providing high-frequency support to guide the human toward the final goal. First, the Manager uses a subtask filter to identify feasible subtasks based on the agent’s and environment’s current states, selecting from the potential subtask list from the overall plan created by the Coordinator. Next, it converts the selected subtasks, coordination plan, current agent states, and environment state into a language-based prompt, querying GPT-4 to generate the next subtask. This query is generated immediately upon completing the current subtask, with an average latency of $\sim$ 1.2 seconds. The determined subtask in every query will then be converted into language feedback and sent to humans.
Greedy Planner
Once given the target subtask, the human determines the atomic actions needed to complete it. For collaborative tasks with an autonomous agent, a greedy planner using Depth First Search (DFS) finds the optimal path with the lowest action cost to finish the subtask.
Multi-Modal feedbacks
Building on the proposed HRT-ML framework, we introduce agents that provide four different language feedbacks described as follows: Inactive Feedback, Passive Feedback, Active Feedback, and Superactive Feedback.
  • Inactive Feedback agent (IFA): The IFA collaborates with humans without language communication and coordination. Only the Manager generates target subtasks for the greedy planner.
  • Passive Feedback agent (PFA): The PFA starts to provide passive feedback only when a human requests. The human player takes the role of the leader, while the agent acts as the follower, passively responding to human requests. If there are no specific human commands, PFA will behave like an IFA.
  • Active Feedback agent (AFA): The AFA collaborates with humans as peers. In this mode, both humans and agents can reach out to give language feedback. For user study, we prompt the GPT-4o coordinator in low frequency ($\sim$ 20s) to analyze human conversation history and suggest coordination strategies.
  • Superactive Feedback agent (SFA): The SFA treats humans as novices, acting as supervisors by providing frequent, continuous guidance on every subtask.
Data collection
In this section, we aim to explore the agent’s ability to provide language feedback to help improve human satisfaction and teaming efficiency. Based on the proposed HRT-ML, we consider four types of language feedback and three different layouts: easy, medium, and difficult, with increasing task difficulty and map complexity (Fig. 3). Human participants will play on each layout paired with an autonomous agent with varying levels of language feedback. Further details will be provided in the following section.
Layouts with different complexities
To test the performance of the four agents and the influence of language feedback, we implemented four overcooked maps (Fig. 3). One of the maps is an introductory map, designed for participants to get familiar with the game and operation. The other three maps have varying levels of difficulty, which we refer to as easy, medium, and hard. The more challenging maps have more “dead-ends”, requiring humans and agents to have a better coordination strategy to maintain the team efficiency. Furthermore, the hard map introduces complexity in task orders by requiring multiple ingredients, which demands that humans and the agent reason about orders containing both tomatoes and onions.
Figure 3: Overview of the human study procedure involving 16 participants, beginning with an introduction map to learn game mechanics and agent behavior followed by trials on Easy, Medium, and Hard Layout with four different agents. Post-session surveys were conducted to collect data on participant satisfaction, engagement, trust, and feedback.
Participants
In this study, we recruited 16 participants (9 male and 7 female) aged between 23 and 33 to evaluate the performance of collaborative agents. We selected participants with varying levels of familiarity with digital agents in game environments: 10 participants with self-reported video game time between 1-10 hours per week, 5 participants reported 10-20 hours per week, and 1 reported 20-30 hours per week. Additionally, 13 of the 16 participants reported prior familiar with Large Language Models (LLMs) and embodied language agents, while 3 of the 16 participants reported no prior knowledge or experiences with language agents.
Procedures
Participants were first asked to complete a consent form per Institutional Review Board (IRB) protocol. Subsequently, prior to the formal trials they were given an opportunity to play with the four agents in an introduction map (Fig. 3) to explore various language feedback styles. During this introduction phase participants were guided through system operations, game rules, and agent functionalities. After the introduction phase, participants proceeded to independently collaborate with each of the four agent types, on easy, medium, and hard layouts (Fig. 3), to complete the maximum score within the given time limit (60 s for each layout). For each participant, the scenario (i.e., agent type $\times$ layout difficulty) was set to show up in a randomized order. We collected a total of 192 experiment trials, 12 trials per participant.
We collected game scores and step-by-step action logs for each experiment shown in (Fig. 3). After completing the teaming scenario, participants were also asked to fill out a survey rating their satisfaction, engagement, and trust level for each experiment, on a seven-point Likert scale. Participants were also asked to specify their preferred language feedback level for each layout.
Additionally, participants were asked to provide improvement suggestions on the language feedback provided by agents (e.g., “If you were to play with this agent in an Overcooked game competition, what changes or improvements would you suggest for the agent’s feedback?”), and state the reason for their satisfaction ratings. For more details, the full questionnaire is available at  [^1] Link to the questionnaire:
https://docs.google.com/forms/d/1xSRPH1mkrZxuKXQ0xvjl9Cv4yi0l-aBblpwyG5MbX34/viewform?edit_requested=true
Results and Discussion
The purpose of our data collection and analysis was to test the following hypotheses: An increased level of language support will result in an increase in the perceived level of trustworthiness and intelligence of the agent, and improve overall team performance.
Surprisingly, our data suggested while language feedback could indeed facilitate human trust, perceived agent intelligence, and team efficiency, the desired level of language support exhibited a different relationship than hypothesized. We report our findings in the subsequent sections.
Language feedback builds human trust and perceived intelligence
Trust is a key factor in human-AI collaboration, shaping human experiences and significantly influencing long-term collaboration efficiency (?). In this section, we report the perceived levels of trust and intelligence after participants teamed with four types of agents across three layouts. Both intelligence and trust levels were measured using a 7-point Likert scale, ranging from “Very Untrustworthy” (or “Very Unintelligent”) to “Very Trustworthy” (or “Very Intelligent”). We found that as the agent’s support level increased from Inactive to Superactive, intelligence rating increased monotonically (Fig. 4). A similar trend was observed in the trust ratings. This suggested that as hypothesized, active communication can facilitate building trust and preceived intelligence. Another interesting observation was that, as the agents become more active, the standard deviation for trust ratings and intelligent ratings also increased. This suggested that both trust and intelligence levels also become more influenced by individual human preferences when the agent takes on a more active role in assisting.
Figure 4: Human perceived agent intelligence level (blue bar) and trust level (red bar) represented on a seven-point Likert scale, ranging from “Very Unintelligent/Untrustworthy” to “Very Intelligent/Trustworthy”.
Appropriate Language feedback improves collaboration efficiency
We used the game score to evaluate the team perfromance and collaboration efficiency between humans and agents. Overall, the team scored more points in easy layout with all agent types (53.6 in average) as compared to the medium (30.6 in average) and hard (27.3 in average) layouts. This is not surprising, as the easy layout has simpler maps and fewer subtasks, requiring minimal coordination to complete. In contrast, medium and hard layouts introduced more complex subtasks and dependencies, requiring greater coordination and team effort, which could increase the chance of errors during the task and result in lower scores.
[…]
Figure 5: Game scores of all participants paired with different agents across various layouts. The score distribution of each agent type, IFA, PFA, AFA, SFA, is represented by the blue, red, green, and yellow boxes, respectively.
However, the team performance was not always better with more active agents. In the easy layout, team with PFA performed best, scoring 83.6 in average, significantly higher than IFA, AFA, and SFA (Fig. 5). It was expected that PFA performed better than IFA in the easy task, as it provided effective support with minimal interference, enabling participants to benefit from its assistance as needed and allowing humans to take the lead. This was also supported by user feedback from the survey – “Robot listening to the user and following the commands would help the game better. In terms of intelligence, passive robots work better but still lack user assistance”. Unexpectedly, while AFA and SFA offered more active feedback, they achieved lower scores than the PFA (Fig. 5), suggesting that constant communication was less effective for simple tasks, and may even distract humans. As the complexity increases to the medium level, the team with SFA exhibited a huge increase in score (Fig. 5), exceeding the performance of PFA, implying that as task complexity increased, the frequent support provided by SFA became more valuable. Similarly, in the hard layout, the SFA and AFA demonstrated superior performance as compared to IFA and PFA (Fig. 5), underscoring the value of active guidance and high-frequency support in facilitating collaboration and helping improve task execution in complex tasks.
Interestingly, despite SFA’s higher frequency of active support compared to AFA, overall performance still decreased due to the high complexity of the hard layout compared to AFA. Participant feedback highlighted this challenge, e.g.,the robot/agent is not as smart as me,” and, “I have to give it a long instruction set, and I am more intelligent than it in the hard layout.” One interpretation of these responses was that participants found the agent’s support insufficient for complex tasks. As a result, instead of reducing cognitive load, the frequent suggestions from the agent required human to carefully think about responses, which ultimately increased cognitive demands and decreased team efficiency. Another interpretation is that psychologically, humans may have the preference to demonstrate and maintain intellectual superiority in challenging tasks when teaming with AI agents. As a result, how AI agents communicate suggestions may greatly influence humans’ acceptance rate and team efficiency.
Overall, our results revealed that, the language feedback provided by LLMs can boost human-robot collaboration efficiency and increase human satisfaction. However, the proactiveness and frequency of the language feedback should be provided based on the task complexity and the capabilities of LLMs.
Humans don’t always prefer the best-performing agents
Interestingly, our data suggested that human does not always prefer the agent type that helped achieve the highest game scores. For example, even though the IFA achieved the lowest scores on easy layout (Fig. 5), over 50% of participants reported IFA as their preferred agent among the four types (Fig. 6A). Similarly, even though the SFA outperformed PFA by almost two folds in terms of team score (Fig. 5), 56% of participants selected the PFA as their preferred agent, and 0% of them preferred SFA (Fig. 6A).
We believe this preference shift can be explained by the flow theory (??), which states that people feel most engaged when task complexity aligns with their skill level, and robots can help achieve this balance by adjusting their level of support. In the easy layout, the task was not significantly beyond human’s skill level, and the need for additional help to maintain engagement was low. Therefore, while the active agents can provide help, this help did not significantly influence human’s engage level and satisfaction rate (Fig. 6B). As a result, the non-active agent, which requires the lowest level of cognitive load (“cost”) was preferred. As the difficulty level increases, however, the gap between task complexity and human skill level increases, resulting a disruption to the task-capability balance and reduced engagement. Meanwhile, the additional feedback and support from the more active agents, such as assigning subtasks (e.g., “pick up the onion from (x, x).”, can reduce the coordination and planning effort on the human, and restore the task-capability balance and enhance human’s feel of engagement and joy. At this point, the cost of communication became negligible as compared to the need for sense of achievement, making more active agents more desirable in these challenging scenarios (Fig. 6B).
Figure 6: Participant preferences and satisfaction levels for different agent types across layouts. (A) Bar chart showing the percentage of agent preference by participants for different layouts. (B) Satisfaction ratings for each agent type in different layouts, represented on a seven-point Likert scale, ranging from “Very Unsatisfied” to “Very Satisfied,” with error bars indicating the standard error of the mean (SEM) in responses.
To test this, we recorded participants’ engagement during each game trial. As shown in Fig. 7, satisfaction levels increased with higher engagement, with a correlation test yielding a p-value of 0.00275 and a correlation coefficient of 0.93, indicating a strong relationship. This strong correlation between satisfaction and engagement indicates that engagement is the cause of different satisfaction levels reported by participants. This aligned with our theory, emphasizing the need to align agent support and feedback frequency based on task complexity relative to human capabilities.
Figure 7: Illustration of the relationship and variability between satisfaction level and the engagement of participants for all experimental trials. The box plot shows how satisfaction levels correspond to participants with an engagement level from 0 (Very distracted) to 6 (Very engaged).
Adaptively assigning subtasks and providing language feedback
Our findings revealed the importance for LLM agents to adapt their language feedback by considering the relationship between task complexity, $T_{h}$, human capability, $C_{h}$, and the LLM’s capability, $C_{l}$. Below we propose a simple adaptation strategy for LLM agents to select their support level and language feedback frequency:
  • $C_{h}>T$ and $C_{l}<T$: Here human capability surpasses the task’s complexity, and the LLM capability is not sufficient to address the challenging tasks without human guidance. Based on our results, a passive (PFA) to relative infrequent (AFA) feedback style would allow the agent to provide sufficient support to improve team performance and request human help when needed, while keeping communication frequency to a minimal to avoid overhead on the human side. This way, human teammates with higher capability level could guide the agent on high-level coordination strategies and specific subtask execution, keeping them engaged.
  • $C_{h}<T$ and $C_{l}>T$: Here task complexity exceeds human capability, while LLM is fully capable of executing the task. In this case, extra active agent feedback and support (SFA) are crucial for maintaining team performance, and help reduce the gap between human capability and task complexity.
  • $C_{h}<T$ and $C_{l}<T$: Here task complexity exceeds human capability. However, LLM capability is also not sufficient to address the challenging task neither. In this case, language feedback from LLM is often not useful in resolving the challenges that human is struggling with, and can even be misleading. High-frequency communications from LLM in this scenario would require additional effort from human to respond, potentially further heightening the anxiety that human is already experiencing (?), decreasing team performance and human engagement. Our results suggested that a more passive (PFA) or relative infrequent (AFA) feedback style would result in better teaming performance in this case.
  • $C_{h}>T$ and $C_{l}>T$: Here both human and LLM capabilities surpass the task’s complexity. The active feedback style (AFA) could allow the human and agents to communicate their needs at a comfortable pace, and improve collaboration efficiency.
Conclusion
In this work, we introduced HRT-ML, a flexible human-robot teaming framework designed to provide adaptive communication feedback to humans at varying levels and frequencies. The HRT-ML framework comprises two core modules: a Coordinator for high-level, low-frequency strategic guidance and a Manager for task-specific, high-frequency instructions, allowing collaborating with humans across four distinct feedback styles: Inactive, Passive, Active, and Superactive.
Our user study results demonstrated that language-based feedback from LLMs can significantly enhance collaboration performance and foster human trust, and that as task complexity increases, more frequent, proactive support is desired. However, our study also revealed that it is critical for the agent to select their language feedback frequency based on task complexity, human capability, and agent capability. Overly frequent feedback in simple tasks or from less capable agents does not effectively increase team performance and satisfaction, and could even increase effort and reduce human engagement. Based on these findings, we proposed a simple principle that allows agents to adapt their language feedback style according to perceived task challenges, human capabilities, and LLM capabilities.
Limitations and future work
In this work, we focused on investigating the effect of agent feedback frequency on team performance. As such, each agent was set to a constant active level, and cannot dynamically adjust their level of support and language feedback throughout the task. In real-world scenarios, task complexity and human skill levels for different sub task can vary dynamically, requiring agents to adjust their communications and behaviors accordingly. Future work should explore methods for adaptive agent to estimate human cognitive load, capabilities, and engagement, and adjust LLM feedback in real time to enable better team performance and adaptability. The results from our study provide the basis for designing and implementing such real-time feedback adjustments. Going forward, these adaptive response and feedback capabilities can empower future LLM agents to flexibly support human needs in a wide variety of task complexities, fostering true partnerships and enhancing teamwork outcomes.
Acknowledgment
This work is supported by the NASA Planetary Science and Technology Through Analog Research (PSTAR) program, Award # 80NSSC22K1313. The authors would like to thank Stefanos Nikolaidis for helpful discussions, and Siddharth Srikanth for assistance with prompt engineering and code refactoring.

### v2

@@
-Effect of Adaptive Communication Support on Human-AI Collaboration
+Effect of Adaptive Communication Support on LLM-powered Human-Robot Collaboration
 Abstract
-Effective human-AI collaboration requires agents to adopt their roles and levels of support based on human needs, task requirements, and complexity.
-Traditional human-AI teaming often relies on a pre-determined robot communication scheme, restricting teamwork adaptability in complex tasks.
+Effective human-robot collaboration requires robot to adopt their roles and levels of support based on human needs, task requirements, and complexity.
+Traditional human-robot teaming often relies on a pre-determined robot communication scheme, restricting teamwork adaptability in complex tasks.
 Leveraging strong communication capabilities of Large Language Models (LLMs), we propose a Human-Robot Teaming Framework with Multi-Modal Language feedback (HRT-ML), a framework designed to enhance human-robot interaction by adjusting the frequency and content of language-based feedback.
-HRT-ML framework includes two core modules: a Coordinator for high-level, low-frequency strategic guidance, and a Manager for task-specific, high-frequency instructions, enabling passive and active interactions with human teammates.
-To assess the impact of language feedback in collaborative scenarios, we conducted experiments in an enhanced Overcooked-AI game environment with varying levels of task complexity (easy, medium, hard) and feedback frequency (inactive, passive, active, superactive).
-Our results show that as task complexity increases relative to human capabilities, human teammates exhibited stronger preference towards robotic agents that can offer frequent, proactive support.
-However, when task complexities exceed the LLM’s capacity, noisy and inaccurate feedback from superactive agents can instead hinder team performance, as it requires human teammates to increase their effort to interpret and respond to the large amount of communications, with limited performance return.
-Our results offers a general principle for robotic agents to dynamically adjust their levels and frequencies of communications to work seamlessly with human and achieve improved teaming performance.
-Introduction
-Human-robot collaboration has been extensively studied and applied across diverse scenarios, demonstrating strong potential towards enhanced efficiency and performance (???????).
-As task complexity increases, agent adaptability becomes increasingly essential for seamless teamwork.
-Previous work has developed methods and tools for robot to adapt their actions based on inferred human objectives (?), trust level (?), and individual preferences (?).
-Recent work also begin to incorporate language-based feedback (??) to enable more direct communications and lower user barriers.
+HRT-ML framework includes two core modules: a Coordinator for high-level, low-frequency strategic guidance, and a Manager for subtask-specific, high-frequency instructions, enabling passive and active interactions with human teammates.
+To assess the impact of language feedback in collaborative scenarios, we conducted experiments in an enhanced Overcooked environment with varying levels of task complexity (easy, medium, hard) and feedback frequency (inactive, passive, active, superactive).
+Our results show that as task complexity increases relative to human capabilities, human teammates exhibited a stronger preference towards robotic agents that can offer frequent, proactive support.
+However, when task complexities exceed the LLM’s capacity, noisy and inaccurate feedback from superactive robotic agents can instead hinder team performance, as it requires human teammates to increase their effort to interpret and respond to a large number of communications, with limited performance return.
+Our results offer a general principle for robotic agents to dynamically adjust their levels and frequencies of communications to work seamlessly with humans and achieve improved teaming performance.
+[…]
+1 Introduction
+Human-robot collaboration has been extensively studied and applied across diverse scenarios, demonstrating strong potential towards enhanced efficiency and performance [20, 10, 29, 15, 35, 24, 25].
+As task complexity increases, robot adaptability becomes increasingly essential for seamless teamwork.
+Previous work has developed methods and tools for robot to adapt their actions based on inferred human objectives [25], trust level [8], and individual preferences [3].
+Recent work also begin to incorporate language-based feedback [28, 31] to enable more direct communications and lower user barriers.
 However, robot communications in these approaches primarily focused on relative simple commands such as “pick up the book” or “move left a little”, which do not reflect the level of human-robot communications required in real-world applications.
-Recent advancements in large language models (LLMs) have brought powerful reasoning (????), natural language understanding (??), contextual awareness (?), and generalization capabilities (?), enabling more advanced, prolonged communications (???).
-These methods has empowered agents to process ambiguous and complex instructions from human (?), engage in more natural and dynamic conversations (??), and learn from a diverse set of inputs (??).
-However, even in these LLM-enhanced communications, human teammates continue to play a predominant role in requesting specific tasks and providing suggestions during collaboration (?).
+Recent advancements in large language models (LLMs) have brought powerful reasoning [36, 38, 1, 16], natural language understanding [23, 34], contextual awareness [12], and generalization capabilities [14], enabling more advanced, prolonged communications [4, 27, 18].
+These methods has empowered robots to process ambiguous and complex instructions from human [23], engage in more natural and dynamic conversations [34, 19], and learn from a diverse set of inputs [14, 32].
+However, even in these LLM-enhanced communications, human teammates continue to play a predominant role in requesting specific tasks and providing suggestions during collaboration [23].
 Schoenegger et al.
-(?)’s study suggested that state-of-the-art LLMs can often match or surpass human performance in various domains.
-Based on these results, we hypothesize that allowing LLM agents to more proactively participate in (?) or even initiate communications with human teammates can enhance teaming performance and efficiency.
+[30]’s study suggested that state-of-the-art LLMs can often match or surpass human performance in various domains.
+Based on these results, we hypothesize that allowing LLM-powered robots to participate more proactively in [33] or even initiate communications with human teammates can enhance teaming performance and efficiency.
 To test this hypothesis and systematically evaluate the impact of different forms of language feedback provided by robots on collaboration efficiency and human satisfaction, in this study we develop HRT-ML, a human-robot teaming framework incorporating multi-modal language feedback to support dynamic, context-aware teaming styles.
@@
 Combining Coordinator and Manager allows the robotic agent to provide different forms of instruction at different frequencies.
-To investigate how the form and frequency of language feedback influence teaming performance, we performed user studies using four different agent active levels: Inactive, Passive, Active, and Superactive.
-We find that as the environment becomes more challenging, participants exhibit stronger preference for agents that provide active support, whereas in simpler tasks, frequent agent feedback is often perceived negatively, reducing overall performance and human satisfaction.
+To investigate how the form and frequency of language feedback influence teaming performance, we performed user studies using four different robot active levels: Inactive, Passive, Active, and Superactive.
+We find that as the environment becomes more challenging, participants exhibit a stronger preference for robots that provide active support, whereas in simpler tasks, frequent feedback is often perceived negatively, reducing overall performance and human satisfaction.
 These findings reveal the importance for the active level of robot language support to dynamically adapt based on task complexity and team capability.
-Inaccurate or overly-frequent feedback can decrease team efficiency, as participants must take extra time to understand and correct these suggestions.
+Inaccurate or overly frequent feedback can decrease team efficiency, as participants must take extra time to understand and correct these suggestions.
 In summary, our contributions are as follows:
   • developed a human-robot teaming framework incorporating multi-modal language feedback to support dynamic, context-aware teaming styles.
-  • performed user studies to determine how forms and frequency of agent language feedback influence teaming performance
-  • discovered how the active level of robot language support should adapt based on task complexity and team capability, to best enhance teaming performance
-To the best of the authors’ knowledge, this is the first study that systematically explores the effect of different types of LLM-based multimodal language feedback on teaming performance across a wide range of task complexities.
-Testbed: Overcooked-AI
-To test the influence of language feedback on human-robot collaboration, we chose Overcooked-AI (?), a platform designed to assess multi-agent coordination skills.
-In this game, players are motivated to collaborate actively to maximize their score by completing orders within a time limit.
-A score of 60 points was awarded when the correct soup is served.
+  • performed user studies to determine how forms and frequency of robot language feedback influence teaming performance.
+  • discovered how the active level of robot language support should adapt based on task complexity and team capability, to best enhance teaming performance.
+To the best of the authors’ knowledge, this is the first study that systematically explores the effect of different types of LLM-based multimodal language feedback on teaming performance across different task complexities.
+2 Related Work
+Robots serving as teammates and collaborating with humans in shared workspaces is an emerging paradigm in human-robot interaction [9, 13, 39].
+Recent advancements in LLMs have further enabled robots to interpret human intent, generate task plans, and engage in more natural interactions [37, 22, 39].
+In our work, we specifically focus on the role of communication in LLM-based human-robot collaboration, as effective information exchange is crucial for task coordination, intent alignment, and adaptive interaction.
+Communication in human-robot collaboration can generally be categorized as either unidirectional or bidirectional.
+In unidirectional communication, robots do not proactively engage with humans; instead, they infer human intent from instructions and make autonomous decisions based on observed cues [36, 6, 22, 23].
+In contrast, bidirectional communication involves robots actively interacting with humans—seeking clarification, sharing updates, and negotiating task execution strategies [16, 37, 26, 39].
+Several studies have explored unidirectional and bidirectional communication strategies, finding that humans perceive robots capable of bidirectional communication as more intelligent and collaborative [2, 39].
+However, existing studies [2, 39] on bidirectional communication primarily focus on updating status or requesting simple, predefined subtasks without incorporating complex reasoning about the coordination strategies.
+A well-defined bidirectional communication framework for complex coordination tasks is still lacking.
+Moreover, the impact of different forms and frequencies of bidirectional communication on coordination performance and human preference remains unexplored.
+3 Testbed: Overcooked-Environment
+To test the influence of language feedback on human-robot collaboration, we chose Overcooked-AI [5], an environment designed to assess multi-agent coordination skills.
+In this game (Figure 1, humans (blue hat agents) and robots (green hat agents) are motivated to collaborate actively to maximize their score by completing orders within a time limit.
+A score of 53 points was awarded when the correct soup was served.
 Partial points will be awarded to incomplete or incorrect soups based on the number of missing/incorrect components.
@@
 1A), and to finish subtasks, chefs need to move and interact with the environment.
-This process can be applied to any collaborative setting: first, reasoning through subtasks to achieve the overall goal, then selecting low-level atomic actions to complete each subtask.
+This decision making process can be applied to any collaborative setting: first, reasoning through subtasks to achieve the overall goal, then selecting low-level atomic actions to complete each subtask.
 Figure 1: (A) Cooking process to complete an order.
 (B) The designed human-AI collaboration interface (left: game layout, right: communication panel).
-The red cross represent an example of the intermediate empty counter used to collaborate.
-Subtasks
-The subtasks in an overcooked environment for a single agent can be broken down into three main parts:
+The Blue hat agent and Green hat agent are operated by humans and robots, respectively.
+The red cross represents an example of the intermediate empty counter used to collaborate.
+3.0.1 Subtasks
+The subtasks in an overcooked environment can be broken down into three main parts:
   1. 1.
@@
   2. 2.
-    Cooking: Once the ingredients are placed in the pot, the agent has to start cooking. A timer on the pot signals when the soup is ready to be served.
+    Cooking: Once the ingredients are placed in the pot, the chef has to start cooking. A timer on the pot signals when the soup is ready to be served.
   3. 3.
     Serving: When the soup is ready, the chefs must collect and clean the dish from the dish dispenser, pour the soup into the dish, and deliver it to the serving location.
-In multi-agent collaborative scenarios, each agent has a different path cost for completing each subtask, such as picking up an onion versus a tomato.
-Some subtasks may be unachievable for certain agents.
-Furthermore, by decomposing a subtask into multiple smaller subtasks that can be performed by multiple agents, the overall time cost can be reduced.
+In collaborative scenarios, the human and robot have a different path cost for completing each subtask, such as picking up an onion versus a tomato.
+Some subtasks may be unachievable for humans or robots.
+Furthermore, by decomposing a subtask into multiple smaller subtasks that can be performed by human and robot together, the overall time cost can be reduced.
 For example, in Fig.
-1, the green agent might place the onion at the red cross point (5,2), allowing the blue agent to pick it up from there and add it to the pot.
-Atomic Action
-To finish a specific subtask, such as picking up an onion, the agent must execute a sequence of atomic actions, including movement commands like up, down, left, right, stay, and interact for picking up or placing objects.
-Figure 2: Flowing Human-Robot Teaming Framework (HRT-ML).
+1, the robot might place the onion at the red cross point (5,2), allowing human to pick it up from there and add it to the pot.
+3.0.2 Atomic Action
+To finish a specific subtask, such as picking up an onion, the chef must execute a sequence of atomic actions, including movement commands like up, down, left, right, stay, and interact for picking up or placing objects.
+Figure 2: Human-Robot Teaming Framework with Multi-Modal Feedback(HRT-ML).
 It contains two modules: the Coordinator and the Manager.
-The Coordinator considers human preferences, formulates overall strategies, and provides low-frequency feedback to guide the collaboration.
-The Manager processes these strategies along with the state information to generate high-frequency subtask instructions for both human players and greedy planners, facilitating efficient task coordination and execution.
-Human-Robot Teaming Framework with Multi-Modal Language Feedback
-To provide adaptive language feedback in human-robot collaboration, HRT-ML includes two core components: the Coordinator and the Manager.
-The Coordinator leads high-level strategy discussions with humans, generating a final coordination plan and offering low-frequency feedback.
-In contrast, the Manager handles detailed subtask allocation and provides high-frequency feedback to guide human players.
-Humans and other agents will receive the feedback and execute low-level actions.
-Coordinator
-The Coordinator is responsible for designing overall collaboration strategies and discussing them with the human partner.
-To support this, we employed a structured chain-of-thought approach to make the suggestions and discussions more effective.
-It begins by retrieving relevant information, including the layout, the overall collaboration goal, rules, and the agent’s state formatted as a text description (Figure 2) using a prompt template, and then queries GPT-4o with it.
-It first analyzes the possible subtasks required to achieve the overall goal, then evaluates the difficulty of each subtask for each agent, and generates a plan that maximizes collaboration efficiency.
-For example, based on the environment (Fig.
-1B), our Coordinator suggests: Given the layout, the human (blue agent) should focus on picking up the tomato and serving the dish, while I will handle the onion and place the dish at location (5, 2) or (6, 2).
-We note that the coordination querying is made conversational, allowing humans to continuously revise the plan through continuous discussion after reviewing the plan.
-During these coordination discussions, the game remains paused until the human chooses to end the conversation.
-The conversation can be initiated proactively by the Coordinator at a low frequency or by the human.
-When a human request is made, we incorporate their suggestions into the coordination plan prompt for initially querying GPT-4, explicitly asking it to evaluate the feasibility of these preferences and propose a corresponding coordination plan.
-Manager
-The Manager assigns subtasks at each stage, providing high-frequency support to guide the human toward the final goal.
-First, the Manager uses a subtask filter to identify feasible subtasks based on the agent’s and environment’s current states, selecting from the potential subtask list from the overall plan created by the Coordinator.
-Next, it converts the selected subtasks, coordination plan, current agent states, and environment state into a language-based prompt, querying GPT-4 to generate the next subtask.
-This query is generated immediately upon completing the current subtask, with an average latency of $\sim$ 1.2 seconds.
-The determined subtask in every query will then be converted into language feedback and sent to humans.
-Greedy Planner
-Once given the target subtask, the human determines the atomic actions needed to complete it.
-For collaborative tasks with an autonomous agent, a greedy planner using Depth First Search (DFS) finds the optimal path with the lowest action cost to finish the subtask.
-Multi-Modal feedbacks
-Building on the proposed HRT-ML framework, we introduce agents that provide four different language feedbacks described as follows: Inactive Feedback, Passive Feedback, Active Feedback, and Superactive Feedback.
+The Coordinator designs overall coordination strategies using Directed Acyclic Graph (DAG) and dynamically add and delete nodes and edges through communication with humans.
+Three types of coordination—graph structure revision, graph node attribute revision, and adding temporary subtasks—are enclosed in a red dashed rectangle.
+A blue subtask node represents tasks that are ready for execution, a yellow node indicates tasks that are not yet ready, and a red node signifies emergency tasks.
+The action cost between two subtasks is indicated with numbers on the edges.
+The Manager interprets the graph-represented coordination strategy to allocate subtasks to humans and robots.
+It provides subtask-level instructions to guide humans toward the overall goal and updates the graph status upon the completion of each subtask.
+4 Human-Robot Teaming Framework with Multi-Modal Language Feedback
+To provide adaptive language feedback in human-robot collaboration, HRT-ML leverages a Directed Acyclic Graph (DAG) to represent subtasks and its dependencies (e.g., finishing an onion soup) for human-robot collaboration.
+Based on the DAG, HRT-ML includes two core modules: the Coordinator and the Manager.
+The Coordinator handles high-level strategy discussions, generating an initial subtask graph and revising it based on conversation.
+The Manager is responsible for allocating detailed subtasks based on the finalized coordination DAG, monitoring the status of assigned subtasks for both the human and the robot, and updating the graph accordingly.
+Once a subtask is determined, a Greedy Action Planner is used to generate atomic actions for the robot.
+4.1 Subtask Graph for Human-Robot Collaboration
+We represent the collaboration goal as a dynamic directed acyclic graph (DAG) (Figure 2) where nodes represent subtasks (e.g., pick onion, cook onion), and edges represent dependencies between tasks.
+Each edge is defined as (parent node, node, cost), where the cost is calculated using the Greedy Action Planner (Section 4.4).
+Each node in the graph contains attributes for different coordination purposes:
+  • Task id: Each task gets a unique id, such that the LLM only needs to return the task id, and we can retrieve all other information based on the task id.
+  • Task Type: Categorized as Putting, Getting, or Operating.
+There exist constraints when assigning different subtasks to robots or humans depending on their state, e.g., a robot that already held an onion can not execute any Getting or Operating tasks.
+  • Status: Represents the execution state, includes ready-to-execute, executing, success, failure, emergency.
+The status of each subtask is considered when an LLM is prompted to assign subtasks for Humans and Robots.
+  • Target Position: Specifies a list of locations in the environment where the task can be performed.
+One subtask can be executed at multiple locations.
+This will enable humans and robots to explore additional collaboration strategies, such as the human placing an onion on an empty counter while the robot retrieves it.
+  • Notes: The subtask node includes a notes attribute that captures human preferences or additional instructions (e.g., the human chef prefers to perform picking up an onion; the human chef is observed to be proficient at picking up onion.).
+  • Parent subtasks: The parent subtasks indicate the prerequisite subtasks that are required to be success before this subtask can be executed.
+  • Priority: Priority is considered when assigning subtasks to humans and robots. It is calculated as the accumulated cost from the current node to the sink (termination) node.
+  • Running time: The running time acts as a counter, recording the duration for which the robot has been assigned to this subtask.
+The subtask graph enables the robot to translate language-based coordination into modifications within the graph, helping the LLM organize conversations and align its knowledge with the human’s.
+4.2 Coordinator
+The Coordinator of the robot is responsible for designing overall collaboration strategies and discussing them with the human partner.
+The initial subtask graph is generated using a one-shot LLM prompt (Appendix 12.1), incorporating recipe and kitchen layout information.
+We then implement a conversation-based graph modification, which can be initiated by the robot at a low frequency with the corresponding prompt (Appendix 12.1) (active mode / superactive mode described in Section 4.5 or proactively by the human through language-based feedback).
+The coordination process can be divided into the following decision tree:
+  • The robot first identifies the type of human request, which falls into one of three categories (Figure 2, red dashed rectangle); if it is unsure, it will further ask a human to clarify.
+    • Subtask Graph Structure Change: The human expresses thoughts on how to complete the overall task.
+This may involve adding nodes, further decomposing a subtask for human and robot execution, or correcting incorrect graph dependencies.
+For example, based on the environment (Figure 1B), one of the possible coordination strategies is that the robot can place the onion at location (5, 2), and the human (blue hat) should focus on picking up onions at (5, 2).
+    • Subtask Attribute Change: The human specifies preferences for subtask assignments.
+For example, based on the environment (Figure 1B), a human says that he would like to handle picking up onions and dishes as well as putting them on the counter (5, 2) and (6, 2).
+    • Temporary Subtask Assignment: The human is not designing an overall collaboration strategy but instead requests the robot to execute a specific temporary subtask.
+For example, if the robot is blocking the human’s path while they attempt to put an onion in the pot, the human may instruct the robot to move to a specific position.
+  • After identifying the type of coordination, the subtask graph is updated using the prompts shown in Appendix 12.1 for each category.
+During these coordination discussions, the game remains paused until the human concludes the conversation.
+4.3 Manager
+The Manager (Figure 2) is responsible for real-time subtask allocation based on the information from the subtask graph and subtask status monitoring.
+4.3.1 Subtask allocation
+The Manager were first asked to analyze the current state of humans and robots along with the list of subtasks that are ready to execute in the subtask graph.
+The prompt (Appendix 12.2) explicitly instructs the robot to consider task priorities and types based on the current state, as well as human-reported preferences for each subtask.
+In addition to subtask assignments, the Manager is also prompted to generate language instructions to guide the human to execute their assigned subtasks toward the final goal.
+This feedback is selectively presented to the human based on the feedback mode described in Section 4.5.
+4.3.2 Subtask status update
+The Manager will be triggered to evaluate the current executing task to update the status given the human and robot action trajectory when an Interact action is executed (Prompt details can be found in Appendix 12.2).
+Once a subtask is identified as success, the subtask nodes without unresolved dependencies will be marked as ready-to-execute, while others remain not-ready until their prerequisites are completed.
+4.4 Greedy Action Planner
+Once a subtask (target location) is assigned, the robot uses a greedy planner based on Depth First Search (DFS) to determine the sequence of atomic actions required to complete the task.
+The planner computes the optimal path with the lowest action cost, considering the current environment layout and chef states.
+4.5 Multi-Modal Feedbacks
+Building on the proposed HRT-ML framework, we introduce robotic agents that provide four different language feedbacks described as follows: Inactive Feedback, Passive Feedback, Active Feedback, and Superactive Feedback.
   • Inactive Feedback Agent (IFA): The IFA collaborates with humans without language communication and coordination. Only the Manager generates target subtasks for the greedy planner.
   • Passive Feedback Agent (PFA): The PFA starts to provide passive feedback only when a human requests.
-The human player takes the role of the leader, while the agent acts as the follower, passively responding to human requests.
+The human player takes the role of the leader, while the robotic agent acts as the follower, passively responding to human requests.
 If there are no specific human commands, PFA will behave like an IFA.
   • Active Feedback Agent (AFA): The AFA collaborates with humans as peers.
-In this mode, both humans and agents can reach out to give language feedback.
+In this mode, both humans and robots can reach out to give language feedback.
 For user study, we prompt the GPT-4o coordinator in low frequency ($\sim$ 20s) to analyze human conversation history and suggest coordination strategies.
   • Superactive Feedback Agent (SFA): The SFA treats humans as novices, acting as supervisors by providing frequent, continuous guidance on every subtask.
-Data collection
-In this section, we aim to explore the agent’s ability to provide language feedback to help improve human satisfaction and teaming efficiency.
-Based on the proposed HRT-ML, we consider four types of language feedback and three different layouts: easy, medium, and difficult, with increasing task difficulty and map complexity (Fig.
+5 Data collection
+In this section, we aim to explore the robot’s ability to provide language feedback to help improve human satisfaction and teaming efficiency.
+Based on the proposed HRT-ML, we consider four types of language feedback and three different layouts: easy, medium, and hard, with increasing task difficulty and map complexity (Fig.
 3).
-Human participants will play on each layout paired with an autonomous agent with varying levels of language feedback.
+Human participants will play on each layout paired with a robotic agent with varying levels of language feedback.
 Further details will be provided in the following section.
-Layouts with different complexities
-To test the performance of the four agents and the influence of language feedback, we implemented four overcooked maps (Fig.
+5.1 Layouts with Different Complexities
+To test the performance of the four robotic agents and the influence of language feedback, we implemented four overcooked maps (Fig.
 3).
@@
 The other three maps have varying levels of difficulty, which we refer to as easy, medium, and hard.
-The more challenging maps have more “dead-ends”, requiring humans and agents to have a better coordination strategy to maintain the team efficiency.
-Furthermore, the hard map introduces complexity in task orders by requiring multiple ingredients, which demands that humans and the agent reason about orders containing both tomatoes and onions.
-Figure 3: Overview of the human study procedure involving 16 participants, beginning with an introduction map to learn game mechanics and agent behavior followed by trials on Easy, Medium, and Hard Layout with four different agents.
+The more challenging maps have more “dead-ends”, requiring humans and robots to have a better coordination strategy to maintain team efficiency.
+Furthermore, the hard map introduces complexity in task orders by requiring multiple ingredients, which demands that humans and the robot reason about orders containing both tomatoes and onions.
+Figure 3: Overview of the human study procedure involving 16 participants, the study begins with a tutorial map to introduce game mechanics.
+Followed by the tutorial, the human players will then play on the easy, medium, and hard layouts with four agents with different language feedback.
 Post-session surveys were conducted to collect data on participant satisfaction, engagement, trust, and feedback.
-Participants
-In this study, we recruited 16 participants (9 male and 7 female) aged between 23 and 33 to evaluate the performance of collaborative agents.
-We selected participants with varying levels of familiarity with digital agents in game environments: 10 participants with self-reported video game time between 1-10 hours per week, 5 participants reported 10-20 hours per week, and 1 reported 20-30 hours per week.
-Additionally, 13 of the 16 participants reported prior familiar with Large Language Models (LLMs) and embodied language agents, while 3 of the 16 participants reported no prior knowledge or experiences with language agents.
-Procedures
-Participants were first asked to complete a consent form per Institutional Review Board (IRB) protocol.
-Subsequently, prior to the formal trials they were given an opportunity to play with the four agents in an introduction map (Fig.
+5.2 Participants
+In this study, we recruited 16 participants (9 male and 7 female) aged between 23 and 33 to evaluate the performance of collaborative robotic agents.
+We selected participants with varying levels of familiarity with digital agents in-game environments: 10 participants self-reported video game time between 1-10 hours per week, 5 participants reported 10-20 hours per week, and 1 reported 20-30 hours per week.
+Additionally, 13 of the 16 participants reported prior familiarity with LLMs and embodied language agents, while 3 of the 16 participants reported no prior knowledge or experiences with language agents.
+5.3 Procedures
+Participants were first asked to complete a consent form per the Institutional Review Board (IRB) protocol.
+Subsequently, prior to the formal trials they were given an opportunity to play with the four robotic agents in an introduction map (Fig.
 3) to explore various language feedback styles.
-During this introduction phase participants were guided through system operations, game rules, and agent functionalities.
-After the introduction phase, participants proceeded to independently collaborate with each of the four agent types, on easy, medium, and hard layouts (Fig.
-3), to complete the maximum score within the given time limit (60 s for each layout).
+During this introduction phase participants were guided through system operations, game rules, and robot functionalities.
+After the introduction phase, participants proceeded to independently collaborate with each of the four agent types on easy, medium, and hard layouts (Fig.
+3) to complete the maximum score within the given time limit (60 s for each layout).
 For each participant, the scenario (i.e., agent type $\times$ layout difficulty) was set to show up in a randomized order.
 We collected a total of 192 experiment trials, 12 trials per participant.
-We collected game scores and step-by-step action logs for each experiment shown in (Fig.
+We collected game scores for each experiment shown in (Fig.
 3).
@@
 Participants were also asked to specify their preferred language feedback level for each layout.
-Additionally, participants were asked to provide improvement suggestions on the language feedback provided by agents (e.g., “If you were to play with this agent in an Overcooked game competition, what changes or improvements would you suggest for the agent’s feedback?”), and state the reason for their satisfaction ratings.
-For more details, the full questionnaire is available at  [^1] Link to the questionnaire:
-https://docs.google.com/forms/d/1xSRPH1mkrZxuKXQ0xvjl9Cv4yi0l-aBblpwyG5MbX34/viewform?edit_requested=true
-Results and Discussion
-The purpose of our data collection and analysis was to test the following hypotheses: An increased level of language support will result in an increase in the perceived level of trustworthiness and intelligence of the agent, and improve overall team performance.
-Surprisingly, our data suggested while language feedback could indeed facilitate human trust, perceived agent intelligence, and team efficiency, the desired level of language support exhibited a different relationship than hypothesized.
+Additionally, participants were asked to provide improvement suggestions on the language feedback provided by robots (e.g., “If you were to play with this robotic agent in an Overcooked game competition, what changes or improvements would you suggest for the agent’s feedback?”), and state the reason for their satisfaction ratings.
+For more details, the full questionnaire is available in the Appendix.
+11.
+6 Results and Discussion
+The purpose of our data collection and analysis was to test the following hypotheses: An increased level of language support will result in an increase in the perceived level of trustworthiness and intelligence of the robot, and improve overall team performance.
+Surprisingly, our data suggested while language feedback could indeed facilitate human trust, perceived robot intelligence, and team efficiency, the desired level of language support exhibited a different relationship than hypothesized.
 We report our findings in the subsequent sections.
-Language feedback builds human trust and perceived intelligence
-Trust is a key factor in human-AI collaboration, shaping human experiences and significantly influencing long-term collaboration efficiency (?).
-In this section, we report the perceived levels of trust and intelligence after participants teamed with four types of agents across three layouts.
+6.1 Language Feedback Builds Human Trust and Perceived Intelligence
+Trust is a key factor in human-AI collaboration, shaping human experiences and significantly influencing long-term collaboration efficiency [8].
+In this section, we report the perceived levels of trust and intelligence after participants teamed with four types of robotic agents across three layouts.
 Both intelligence and trust levels were measured using a 7-point Likert scale, ranging from “Very Untrustworthy” (or “Very Unintelligent”) to “Very Trustworthy” (or “Very Intelligent”).
-We found that as the agent’s support level increased from Inactive to Superactive, intelligence rating increased monotonically (Fig.
+We found that as the robot’s support level increased from Inactive to Superactive, intelligence rating increased monotonically (Fig.
 4).
@@
 This suggested that as hypothesized, active communication can facilitate building trust and preceived intelligence.
-Another interesting observation was that, as the agents become more active, the standard deviation for trust ratings and intelligent ratings also increased.
-This suggested that both trust and intelligence levels also become more influenced by individual human preferences when the agent takes on a more active role in assisting.
-Figure 4: Human perceived agent intelligence level (blue bar) and trust level (red bar) represented on a seven-point Likert scale, ranging from “Very Unintelligent/Untrustworthy” to “Very Intelligent/Trustworthy”.
-Appropriate Language feedback improves collaboration efficiency
-We used the game score to evaluate the team perfromance and collaboration efficiency between humans and agents.
-Overall, the team scored more points in easy layout with all agent types (53.6 in average) as compared to the medium (30.6 in average) and hard (27.3 in average) layouts.
+Another interesting observation was that, as the robot become more active, the standard deviation for trust ratings and intelligent ratings also increased.
+This suggested that both trust and intelligence levels also become more influenced by individual human preferences when the robot takes on a more active role in assisting.
+Figure 4: Human perceived robot intelligence level (blue bar) and trust level (red bar) represented on a seven-point Likert scale, ranging from “Very Unintelligent/Untrustworthy” to “Very Intelligent/Trustworthy”.
+6.2 Appropriate Language Feedback Improves Collaboration Efficiency
+We used the game score to evaluate the team performance and collaboration efficiency between humans and robots.
+Overall, the team scored more points in the easy layout with all agent types (53.6 on average) as compared to the medium (30.6 on average) and hard (27.3 on average) layouts.
 This is not surprising, as the easy layout has simpler maps and fewer subtasks, requiring minimal coordination to complete.
@@
 […]
-Figure 5: Game scores of all participants paired with different agents across various layouts.
+Figure 5: Game scores of all participants paired with different robotic agents across various layouts.
 The score distribution of each agent type, IFA, PFA, AFA, SFA, is represented by the blue, red, green, and yellow boxes, respectively.
-However, the team performance was not always better with more active agents.
+However, the team performance was not always better with more active robotic agents.
 In the easy layout, team with PFA performed best, scoring 83.6 in average, significantly higher than IFA, AFA, and SFA (Fig.
@@
 Interestingly, despite SFA’s higher frequency of active support compared to AFA, overall performance still decreased due to the high complexity of the hard layout compared to AFA.
-Participant feedback highlighted this challenge, e.g.,the robot/agent is not as smart as me,” and, “I have to give it a long instruction set, and I am more intelligent than it in the hard layout.” One interpretation of these responses was that participants found the agent’s support insufficient for complex tasks.
-As a result, instead of reducing cognitive load, the frequent suggestions from the agent required human to carefully think about responses, which ultimately increased cognitive demands and decreased team efficiency.
-Another interpretation is that psychologically, humans may have the preference to demonstrate and maintain intellectual superiority in challenging tasks when teaming with AI agents.
-As a result, how AI agents communicate suggestions may greatly influence humans’ acceptance rate and team efficiency.
+Participant feedback highlighted this challenge, e.g.,the robot/agent is not as smart as me,” and, “I have to give it a long instruction set, and I am more intelligent than it in the hard layout.” One interpretation of these responses was that participants found the robot’s support insufficient for complex tasks.
+As a result, instead of reducing cognitive load, the frequent suggestions from the robot required human to carefully think about responses, which ultimately increased cognitive demands and decreased team efficiency.
+Another interpretation is that psychologically, humans may have the preference to demonstrate and maintain intellectual superiority in challenging tasks when teaming with robots.
+As a result, how robotic agents communicate suggestions may greatly influence humans’ acceptance rate and team efficiency.
 Overall, our results revealed that, the language feedback provided by LLMs can boost human-robot collaboration efficiency and increase human satisfaction.
 However, the proactiveness and frequency of the language feedback should be provided based on the task complexity and the capabilities of LLMs.
-Humans don’t always prefer the best-performing agents
+6.3 Humans Don’t Always Prefer the Best-performing Robotic Agents
 Interestingly, our data suggested that human does not always prefer the agent type that helped achieve the highest game scores.
@@
 6A).
-We believe this preference shift can be explained by the flow theory (??), which states that people feel most engaged when task complexity aligns with their skill level, and robots can help achieve this balance by adjusting their level of support.
+We believe this preference shift can be explained by the flow theory [11, 7], which states that people feel most engaged when task complexity aligns with their skill level, and robots can help achieve this balance by adjusting their level of support.
 In the easy layout, the task was not significantly beyond human’s skill level, and the need for additional help to maintain engagement was low.
-Therefore, while the active agents can provide help, this help did not significantly influence human’s engage level and satisfaction rate (Fig.
+Therefore, while the active robotic agents can provide help, this help did not significantly influence human’s engage level and satisfaction rate (Fig.
 6B).
@@
 As the difficulty level increases, however, the gap between task complexity and human skill level increases, resulting a disruption to the task-capability balance and reduced engagement.
-Meanwhile, the additional feedback and support from the more active agents, such as assigning subtasks (e.g., “pick up the onion from (x, x).”, can reduce the coordination and planning effort on the human, and restore the task-capability balance and enhance human’s feel of engagement and joy.
-At this point, the cost of communication became negligible as compared to the need for sense of achievement, making more active agents more desirable in these challenging scenarios (Fig.
+Meanwhile, the additional feedback and support from the more active robotic agents, such as assigning subtasks (e.g., “pick up the onion from (x, x).”, can reduce the coordination and planning effort on the human, and restore the task-capability balance and enhance human’s feel of engagement and joy.
+At this point, the cost of communication became negligible as compared to the need for sense of achievement, making more active robotic agents more desirable in these challenging scenarios (Fig.
 6B).
@@
 This strong correlation between satisfaction and engagement indicates that engagement is the cause of different satisfaction levels reported by participants.
-This aligned with our theory, emphasizing the need to align agent support and feedback frequency based on task complexity relative to human capabilities.
+This aligned with our theory, emphasizing the need to align robot support and feedback frequency based on task complexity relative to human capabilities.
 Figure 7: Illustration of the relationship and variability between satisfaction level and the engagement of participants for all experimental trials.
 The box plot shows how satisfaction levels correspond to participants with an engagement level from 0 (Very distracted) to 6 (Very engaged).
-Adaptively assigning subtasks and providing language feedback
-Our findings revealed the importance for LLM agents to adapt their language feedback by considering the relationship between task complexity, $T_{h}$, human capability, $C_{h}$, and the LLM’s capability, $C_{l}$.
-Below we propose a simple adaptation strategy for LLM agents to select their support level and language feedback frequency:
+7 Adaptively assigning subtasks and providing language feedback
+Our findings revealed the importance of LLM-powered robots adapting their language feedback by considering the relationship between task complexity, $T_{h}$, human capability, $C_{h}$, and the LLM’s capability, $C_{l}$.
+Below we propose a simple adaptation strategy for LLM-powered robots to select their support level and language feedback frequency:
   • $C_{h}>T$ and $C_{l}<T$: Here human capability surpasses the task’s complexity, and the LLM capability is not sufficient to address the challenging tasks without human guidance.
-Based on our results, a passive (PFA) to relative infrequent (AFA) feedback style would allow the agent to provide sufficient support to improve team performance and request human help when needed, while keeping communication frequency to a minimal to avoid overhead on the human side.
-This way, human teammates with higher capability level could guide the agent on high-level coordination strategies and specific subtask execution, keeping them engaged.
+Based on our results, a passive (PFA) to relative infrequent (AFA) feedback style would allow the robot to provide sufficient support to improve team performance and request human help when needed, while keeping communication frequency to a minimal to avoid overhead on the human side.
+This way, human teammates with higher capability level could guide the robot on high-level coordination strategies and specific subtask execution, keeping them engaged.
   • $C_{h}<T$ and $C_{l}>T$: Here task complexity exceeds human capability, while LLM is fully capable of executing the task.
-In this case, extra active agent feedback and support (SFA) are crucial for maintaining team performance, and help reduce the gap between human capability and task complexity.
+In this case, extra active robot feedback and support (SFA) are crucial for maintaining team performance, and help reduce the gap between human capability and task complexity.
   • $C_{h}<T$ and $C_{l}<T$: Here task complexity exceeds human capability.
@@
 In this case, language feedback from LLM is often not useful in resolving the challenges that human is struggling with, and can even be misleading.
-High-frequency communications from LLM in this scenario would require additional effort from human to respond, potentially further heightening the anxiety that human is already experiencing (?), decreasing team performance and human engagement.
+High-frequency communications from LLM in this scenario would require additional effort from human to respond, potentially further heightening the anxiety that human is already experiencing [21], decreasing team performance and human engagement.
 Our results suggested that a more passive (PFA) or relative infrequent (AFA) feedback style would result in better teaming performance in this case.
   • $C_{h}>T$ and $C_{l}>T$: Here both human and LLM capabilities surpass the task’s complexity.
-The active feedback style (AFA) could allow the human and agents to communicate their needs at a comfortable pace, and improve collaboration efficiency.
-Conclusion
+The active feedback style (AFA) could allow the human and robots to communicate their needs at a comfortable pace, and improve collaboration efficiency.
+8 Conclusion
 In this work, we introduced HRT-ML, a flexible human-robot teaming framework designed to provide adaptive communication feedback to humans at varying levels and frequencies.
@@
 Our user study results demonstrated that language-based feedback from LLMs can significantly enhance collaboration performance and foster human trust, and that as task complexity increases, more frequent, proactive support is desired.
-However, our study also revealed that it is critical for the agent to select their language feedback frequency based on task complexity, human capability, and agent capability.
-Overly frequent feedback in simple tasks or from less capable agents does not effectively increase team performance and satisfaction, and could even increase effort and reduce human engagement.
-Based on these findings, we proposed a simple principle that allows agents to adapt their language feedback style according to perceived task challenges, human capabilities, and LLM capabilities.
-Limitations and future work
-In this work, we focused on investigating the effect of agent feedback frequency on team performance.
-As such, each agent was set to a constant active level, and cannot dynamically adjust their level of support and language feedback throughout the task.
-In real-world scenarios, task complexity and human skill levels for different sub task can vary dynamically, requiring agents to adjust their communications and behaviors accordingly.
-Future work should explore methods for adaptive agent to estimate human cognitive load, capabilities, and engagement, and adjust LLM feedback in real time to enable better team performance and adaptability.
+However, our study also revealed that it is critical for the robot to select their language feedback frequency based on task complexity, human capability, and robot capability.
+Overly frequent feedback in simple tasks or from less capable robots does not effectively increase team performance and satisfaction, and could even increase effort and reduce human engagement.
+Based on these findings, we proposed a simple principle that allows robots to adapt their language feedback style according to perceived task challenges, human capabilities, and LLM capabilities.
+9 Limitations
+In this work, we focused on investigating the effect of robot feedback frequency on team performance.
+As such, robots were set to a constant active level and could not dynamically adjust their level of support and language feedback throughout the task.
+In real-world scenarios, task complexity and human skill levels for different sub-tasks can vary dynamically, requiring robots to adjust their communications and behaviors accordingly.
+Future work should explore methods for the adaptive robots to estimate human cognitive load, capabilities, and engagement and adjust LLM feedback in real-time to enable better team performance and adaptability.
 The results from our study provide the basis for designing and implementing such real-time feedback adjustments.
-Going forward, these adaptive response and feedback capabilities can empower future LLM agents to flexibly support human needs in a wide variety of task complexities, fostering true partnerships and enhancing teamwork outcomes.
+Going forward, these adaptive response and feedback capabilities can empower future LLM-powered robots to flexibly support human needs in a wide variety of task complexities, fostering true partnerships and enhancing teamwork outcomes.
 Acknowledgment
-This work is supported by the NASA Planetary Science and Technology Through Analog Research (PSTAR) program, Award # 80NSSC22K1313.
-The authors would like to thank Stefanos Nikolaidis for helpful discussions, and Siddharth Srikanth for assistance with prompt engineering and code refactoring.
+We omitted the acknowledgment to comply with the double-blind review policy.